# Phonon Induced Energy Relaxation in Quantum Critical Metals

### v1

@@ Abstract
Proceeding as before, and assuming $\varphi$ well equilibrated with the electrons at $T_{e}$, with $T_{e}\approx T_{p}\approx T$, we obtain $\partial_{t}E=\kappa^{(2)}(T_{e}-T_{p})$, where
$\kappa^{(2)}=\frac{1}{4}\sum_{i}\int\frac{{\rm d}^{3}\bm{q}}{(2\pi)^{3}}\left|N^{i}_{\bm{q}}\right|^{2}\left(-A_{\varphi}(c{\left|\bm{q}\right|},\bm{q})\right)\frac{c{\left|\bm{q}\right|}}{2T^{2}\sinh^{2}\frac{c{\left|\bm{q}\right|}}{2T}}\,.$ (8)
Compared to Eq. (5), the contribution from the particle-hole bubble is replaced by the $\varphi$-boson spectral function. To analyze the integral for $T\ll{\omega_{\rm D}}$, we make the following simplifications: (1) The typical phonon momentum ${\left|\bm{q}\right|}$ is replaced by $T/c$. (2) We approximate the matrix element by Goldstone’s theorem $\sum_{i}|N_{\bm{q}}^{i}|^{2}\to A_{N}{\left|\bm{q_{\text{2D}}}\right|}^{2}$, and we return to the neglected $q_{z}$ component later. With these, the integral simplifies to
@@
$\kappa^{(2)}\approx\frac{A_{N}}{c^{5}}\int_{0}^{\pi}{\rm d}\theta\frac{\gamma T^{5}\sin^{3}\theta}{\gamma^{2}T^{2}+\left[m^{2}(T)+(\sin^{2}\theta)T^{2}v_{\varphi}^{2}/c^{2}\right]^{2}}\,.$ (9)
Here $\theta$ is the angle between phonon momentum $\bm{q}$ and the $z$-axis [79].
@@
$\kappa^{(2)}$ exhibits a complex sequence of crossovers marked by different power-laws as a function of temperature, as illustrated in Fig. 2 (b) and Fig. 3(a). They are as follows:
@@
(i) Let us start slightly away from the QCP, $m(T=0)=m_{0}>0$, which is assumed to be smaller than all other energy scales, including ${\omega_{\rm D}}$. At the lowest temperature, the denominator of Eq. (9) is dominated by $m_{0}$, so we obtain $\kappa^{(2)}\sim(A_{N}/c^{5})(\gamma/m_{0}^{4})T^{5}$. In this regime, $\varphi$ mediates local interaction between the electrons, and the effects of damping and their dispersion are suppressed relative to $m_{0}$. Alternatively, by counting the phonon phase-space, energy transfer and expansion around the thermal equilibrium, we can obtain $\kappa^{(2)}\sim A_{N}T^{4}\times\braket{\varphi\varphi}$. In this low-energy and long-wavelength limit, the boson $\varphi$ is expected to behave roughly as an ohmic bath, which yields $\braket{\varphi\varphi}\propto\gamma T/m_{0}^{4}$. We also remark in passing that these bosonic collective modes are generically present in interacting Fermi liquids, and $\kappa^{(2)}$ can be interpreted as an interaction correction to energy relaxation on top of the more dominant free fermion result $\kappa^{(1)}$. When the system is deep in the Fermi liquid phase, the boson mass $m_{0}$ is expected to be an electronic energy scale that overwhelms other crossovers, i.e. $\kappa^{(2)}$ grows as $T^{5}$ until being cutoff at $T\sim{\omega_{\rm D}}$.
@@
(iii) When $T\gtrsim\gamma(c^{2}/v_{\varphi}^{2})$, the $T^{2}v_{\varphi}^{2}/c^{2}$ term in the denominator becomes relevant. The $\theta$-integral in Eq. (9) should be cutoff at $\theta_{\text{min}}\sim(c/v_{\varphi})\Delta(T)/T$, which yields $\kappa^{(2)}\sim A_{N}/(cv_{\varphi}^{4})\gamma T\ln(v_{\varphi}^{2}T/(c^{2}\gamma))$, which signals an enhanced contribution due to phonons moving in the $z$ direction. In this regime, the typical phonon momentum is large compared to the typical $\varphi$ frequencies, and unlike the previous lower-temperature regime, the phonons are now sensitive to diffusive character of $\varphi$. This regime is unique to phonons for the following reason. Suppose if we were to replace the phonons by a distinct bosonic mode of electronic origin, but now with $c\approx v_{\varphi}$. Then this regime would onset at $T\sim\gamma c^{2}/v_{\varphi}^{2}\approx\gamma$, which is a high electronic energy scale. In the present setting associated with phonons, the onset of the $\kappa^{(2)}\sim T\ln(T)$ regime is tied to the parametric suppression in $c\ll v_{\varphi}$.
@@
It is useful to make a few remarks about the underlying dimensionality of the problem. The lowest temperature regimes (i), (ii) exhibiting $\kappa^{(2)}\sim T^{5}$ and $\kappa^{(2)}\sim T^{3}$ depend on the phonon dimensionality in a straightforward fashion via their density of states. However, the $\kappa^{(2)}\sim T\ln(T)$ regime (iii) involves a subtle kinematic interplay between the phonon and $\varphi$. In particular, the phonon dispersion along the $z-$direction contributes more significantly compared to the other directions, leading to the $\ln$ enhancement; the latter would be absent if the phonons were purely two-dimensional. Furthermore, $\kappa^{(2)}$ in this regime also depends crucially on the assumption of completely decoupled nature of the two-dimensional electronic layers, via the simplification, $N^{i}_{\bm{q}}\propto{\left|\bm{q_{\text{2D}}}\right|}$, instead ${\left|\bm{q}\right|}$. Finally, let us also consider the additional perturbation due to coupling the $\varphi$ boson to out-of-plane strain, which leads to a correction, $\lambda q_{z}^{2}$ term in $|N^{i}_{\bm{q}}|^{2}$ with $\lambda\ll 1$. This does not lead to qualitatively different results in regimes (i) and (ii), but it can enhance the $\ln$ factor in regime (iii) to a linear factor of $v_{\varphi}^{2}T/(c^{2}\gamma)$ [79]. Since this regime is upper bounded by $T\sim\omega_{D}$, this term remains a small perturbation if $\lambda\ll\gamma c^{2}/(\omega_{D}v_{\varphi}^{2})$.
@@
Energy-relaxation due to non-linear coupling between phonon and collective mode.- Finally, we consider the contribution due to ${\mathcal{L}}_{ep}^{(3)}$ in Eq. (4c), with a phonon self-energy shown in Fig. 1(d), which involves scattering between $\varphi^{2}$ and the acoustic phonons. The renormalized spectral function for $\varphi$ is still given by Eq. (7). The general expression for the flux $\kappa^{(3)}$ reads
@@
Once again, there is a complex sequence of temperature dependent crossovers that can be obtained on the basis of following approximations [79]: replace $\sum_{i}|L^{i}_{\bm{k}+\bm{q}/2,\bm{q}}|^{2}$ by $A_{L}{\left|\bm{q_{\text{2D}}}\right|}^{2}$, and for $T\lesssim{\omega_{\rm D}}$, include contributions from phase-space where $c{\left|\bm{q}\right|}\sim\nu\sim T$. Then $\kappa^{(3)}$ is schematically given by,
$\begin{split}\kappa^{(3)}&\approx\frac{A_{L}}{a_{z}c^{5}}T^{5}\int_{0}^{\pi}{\rm d}\theta\sin^{3}\theta\int{\rm d}^{2}\bm{k}_{\text{2D}}\\ &\times A_{\varphi}(T,\bm{k}_{\text{2D}})A_{\varphi}(T,\bm{k}_{\text{2D}}+\bm{q}_{\text{2D}}(\theta))\,,\end{split}$ (11)
@@
The regimes where $\kappa^{(3)}$ exhibits the distinct $T-$dependent crossovers are identical to the ones that arise in the discussion for $\kappa^{(2)}$ above. (i) For $T\lesssim m_{0}^{2}/\gamma$, the $\varphi$-spectral functions are dominated by $m_{0}$ and the $\bm{k}$-integral can be estimated as $A_{\varphi}^{2}(T,k_{\text{typ}})k_{\text{typ}}^{2}$ treating the $\bm{q}_{\text{2D}}(\theta)$ term as being unimportant. Here $k_{\text{typ}}$ is the typical magnitude of $\bm{k}$, estimated to be $k_{\text{typ}}\sim m_{0}/v_{\varphi}$. Therefore, we obtain $\kappa^{(3)}\sim(A_{L}/a_{z}c^{5}v_{\varphi}^{2})\gamma^{2}T^{7}/m_{0}^{6}$, which is strongly suppressed compared to $\kappa^{(2)}$, as one might expect based on scaling arguments. Alternatively, a simple counting argument leads to $\kappa^{(3)}\sim A_{L}T^{4}\times\braket{\varphi^{4}}$. Here $\braket{\varphi^{4}}\propto T^{3}/m_{0}^{6}$ where the $T^{2}/m_{0}^{8}$ comes from two ohmic baths and there are $N\propto Tm_{0}^{2}$ different ways to distribute the energy between and the momentum these two baths. (ii) As before, when $m_{0}^{2}/\gamma<T<\gamma(c^{2}/v_{\varphi}^{2})$, we can replace $m_{0}^{2}$ by $\gamma T$, and obtain $\kappa^{(3)}\sim(A_{L}/a_{z}c^{5}v_{\varphi}^{2})T^{4}/\gamma$. If we take $m_{0}\rightarrow 0$ in the quantum critical region, these regimes crossover smoothly into each other with $\kappa^{(3)}\sim(A_{L}/a_{z}c^{5}v_{\varphi}^{2})\gamma^{2}T^{7}/m^{6}(T)$.
(iii) When $\gamma(c^{2}/v_{\varphi}^{2})<T<{\omega_{\rm D}}$, the $\bm{q}_{\text{2D}}(\theta)$ term in the second $A_{\varphi}(...)$ factor becomes more important than $\bm{k}_{\text{2D}}$, leading to $\kappa^{(3)}\sim(A_{L}/a_{z}cv_{\varphi}^{6})\gamma T^{2}\ln(v_{\varphi}^{2}T/(c^{2}\gamma))$. The logarithmic factor has the same origin as that in $\kappa^{(2)}$. In the quantum critical regime, this becomes $\kappa^{(3)}\sim(A_{L}/a_{z}cv_{\varphi}^{6})(\gamma^{2}T^{3}/m^{2}(T))\ln(v_{\varphi}^{2}T/(c^{2}\gamma))$. Finally, when $T>{\omega_{\rm D}}$, $\kappa^{(3)}\sim(A_{L}/a_{z}cv_{\varphi}^{6})\gamma T{\omega_{\rm D}}\ln(v_{\varphi}^{2}{\omega_{\rm D}}/(c^{2}\gamma))$. Here, the high-$T$ limit of $\kappa^{(3)}$ is different from $\kappa^{(1)}$ and $\kappa^{(2)}$ in that it does not saturate but grows linearly in $T$. This is expected as $\kappa^{(3)}$ involves an additional boson compared to $\kappa^{(2)}$, and therefore the result should be proportional to the density of that additional boson, which is proportional to $T$, and this will be ultimately cut off at an electronic energy scale where $\varphi$ reveals its electronic nature. A comparison between analytical estimate and numerical result is shown in Fig. 3(b).
@@
Connection to Experiments.- Recent optical experiments involving the technique of two-dimensional coherent spectroscopy [82, 83, 84] have studied the nonlinear response of correlated metals and its relationship to energy relaxation [85, 39], providing a further impetus to a careful theoretical examination of the questions raised above. Over a broad range of hole-dopings in a specific cuprate family [86, 87], recent pump-probe measurements [39] have identified energy relaxation as the dominant nonlinear process that governs the experimental phenomenology. At low temperatures across most of the doping phase diagram, $\Gamma_{E}\sim k_{B}T/\hbar$, with a rate that is significantly slower (approximately by an order of magnitude) compared to the Planckian momentum-relaxation rate [1, 88]. With increasing temperature, there is a tendency for $\Gamma_{E}$ to saturate across a doping dependent crossover scale. Both of these regimes are in stark contrast to the results expected in a conventional Fermi liquid coupled to acoustic phonons [31]. In THz experiments, it is the second rate $\Gamma_{ep}$ that is measured, which leads to a fast initial decay of the pump-probe signal to a plateau at the order of pico seconds, whereas the phonon-environment relaxation $\Gamma_{p-env}$ leads to a slower decay on longer timescales.
@@
To make contact with experiments, we can convert $\kappa=[\kappa^{(1)}+\kappa^{(2)}+\kappa^{(3)}]$ to the energy relaxation rate, $\Gamma_{E}$, via Eq. (2). Here, due to the comparison between the heat capacities $C_{p}\sim(T/{\omega_{\rm D}})^{3}$ and $C_{e}\sim(T/E_{F})\ln(\gamma/T)$, an additional crossover is present, whose temperature scale $T_{*}\sim\sqrt{{\omega_{\rm D}}^{2}/E_{F}}$ is set by equating $C_{p}$ and $C_{e}$. Eq. (2) states that the sector with the smaller heat capacity dominates as its temperature changes more rapidly. When $T\ll T_{*}$, we have $\Gamma_{E}\approx\kappa/C_{p}$ and when $T\gg T_{*}$, $\Gamma_{E}\approx\kappa/C_{e}$. From thermodynamic measurements in a representative family of the cuprates [89], $T_{*}$ is at the order of $1$-$10$ K. For the experimental range of $T\sim O(100\textrm{ K})$, the latter is more relevant, so we would then divide the scaling of $\kappa$ by a power of $T$ to get scaling of $\Gamma_{E}$. Our theoretical results for $\kappa^{(3)}$ provide an appealing scenario for the experimental observations. In the experimental temperature range, the energy relaxation rate $\Gamma_{E}$ is related to $\kappa^{(3)}$ by $\Gamma_{E}\approx\kappa^{(3)}/C_{e}$. Therefore, in regime (iii) and regime (iv) we would obtain a crossover from $T\ln T$ to $T$-independent constant in $\Gamma_{E}$, which qualitatively agrees with the observation in [39] for samples near optimal doping (see Fig. 6c in [39]).
We also comment briefly on the momentum-drag phenomenon [90, 91, 17, 19], which plays a potentially important role in electrical transport in clean systems. In a translational invariant system, the momentum injected into the system is passed around rapidly between the electronic and (critical) bosonic subsystems via rapid collisions, while the system as a whole conserves total momentum. Electrical transport is then determined by the dissipation rate of the total momentum. However, the energy relaxation problem we study in this work is different, since we assume that the phenomenology is dominated by how fast the energy is passed from the hot electron sector to the cold phonon sector, and the dissipation of the total energy to the environment happens at a longer time scale. Therefore, if we model the electron and the phonon as an isolated system, the physics of interest happens before the time scale of drag effects.
Outlook.- We have demonstrated here that in the vicinity of an electronic quantum critical point, the symmetry-allowed coupling between the electronic degrees of freedom and acoustic phonons can lead to a set of complex temperature-dependent crossovers in the energy relaxation rate. When the momentum relaxation in a marginal Fermi liquid is controlled primarily by electronic mechanisms and its interplay with disorder, while the energy relaxation is limited by the flux to the phononic bath, there is no a priori reason for the two relaxation rates to exhibit a similar behavior. Indeed, our theoretical results have unambiguously demonstrated that energy relaxation proceeds via a multitude of crossovers and temperature dependencies across emergent low-energy scales that are distinct from any momentum relaxation rate.
In recent years, inspired in part by experiments [92, 93, 94], there have been suggestions that even the phonons in a strongly correlated metal can lose their quasiparticle-like character [90, 95, 96, 97, 91]. The common underlying theme is that the phonon lifetime is also Planckian, $O(\hbar/k_{B}T)$. Our theoretical conclusions for the energy flux should broadly remain applicable to these incoherent phonons as well in spite of broadened phonon “quasiparticle-peak”, since their typical frequency is still of order $k_{B}T$. A detailed theoretical study of an entangled liquid made out of a non-Fermi liquid of electrons and phonons is an interesting problem for the future.
@@
In this work, while we have focused only on the normal state and ignored the low-temperature superconducting phase, it is likely that the pairing fluctuations above superconducting $T_{c}$ as well as the correlated Bogoliubov quasiparticles below $T_{c}$ , will affect the energy relaxation processes via their coupling to phonons [98], which remains an exciting topic for future research. Finally, given the complex set of crossovers in the energy-relaxation demonstrated by our results near the quantum critical point due to the mismatch between the velocities of the phonons and collective modes, and their dimensionality, we anticipate similar regimes to also appear in the sound attenuation [99, 100, 101, 102, 103, 104, 34]. We leave a detailed analysis of sound attenuation in quantum critical metals for future investigation.

### v2

@@ Abstract
Assuming the $\varphi$ boson is equilibrated with the electrons at $T_{e}\approx T_{p}\equiv T$, the energy transfer rate takes the form $\partial_{t}E=\kappa^{(2)}(T_{e}-T_{p})$, where
$\kappa^{(2)}=\frac{1}{4}\sum_{i}\int\frac{d^{3}{\bm{q}}}{(2\pi)^{3}}|N^{i}_{{\bm{q}}}|^{2}\,[-A_{\varphi}(c|{\bm{q}}|,{\bm{q}})]\,\frac{c|{\bm{q}}|}{2T^{2}\sinh^{2}(c|{\bm{q}}|/2T)}.$ (7)
For $T\ll\omega_{D}$, we approximate $|{\bm{q}}|\sim T/c$ and, using Goldstone’s theorem, take $\sum_{i}|N^{i}_{{\bm{q}}}|^{2}\approx A_{N}{\left|{\bm{q_{\text{2D}}}}\right|}^{2}$. This yields
@@
$\kappa^{(2)}\approx\frac{A_{N}}{c^{5}}\int_{0}^{\pi}d\theta\,\frac{\gamma T^{5}\sin^{3}\theta}{\gamma^{2}T^{2}+\left[m^{2}(T)+T^{2}v_{\varphi}^{2}\sin^{2}\theta/c^{2}\right]^{2}},$ (8)
where $\theta$ is the angle between $\bm{q}$} and the $z$-axis.
@@
$\kappa^{(2)}$ exhibits a complex sequence of crossovers marked by different power-laws as a function of temperature, as illustrated in Fig. 2 and Fig. 3 (a). They are as follows:
@@
(i) Let us start slightly away from the QCP, $m(T=0)=m_{0}>0$, which is assumed to be smaller than all other energy scales, including $\omega_{\rm D}$}. At the lowest temperature, the denominator of Eq. (8) is dominated by $m_{0}$, so we obtain $\kappa^{(2)}\sim(A_{N}/c^{5})(\gamma/m_{0}^{4})T^{5}$. In this regime, $\varphi$ mediates local interaction between the electrons, and the effects of damping and dispersion are suppressed relative to $m_{0}$. Alternatively, the result can be derived from $\kappa^{(2)}\sim A_{N}T^{4}\times\braket{\varphi\varphi}$, where the $T^{4}$ prefactor is the result of phonon phase counting and Gold-stone coupling. In the low-energy and long-wavelength limit, the boson $\varphi$ is expected to behave as an ohmic bath, which yields $\braket{\varphi\varphi}\propto\gamma T/m_{0}^{4}$. We also remark in passing that similar bosonic collective modes are generically present in interacting Fermi liquids, and $\kappa^{(2)}$ can be interpreted as an interaction correction to energy relaxation on top of the more dominant free fermion result $\kappa^{(1)}$. When the system is deep in the Fermi liquid phase, the boson mass $m_{0}$ is expected to be an electronic energy scale that overwhelms other crossovers, i.e. $\kappa^{(2)}$ grows as $T^{5}$ until being cutoff at $T\sim{\omega_{\rm D}}$.
@@
(iii) When $T\gtrsim\gamma(c^{2}/v_{\varphi}^{2})$, the $T^{2}v_{\varphi}^{2}/c^{2}$ term in the denominator becomes relevant. The $\theta$-integral in Eq. (8) should be cutoff at $\theta_{\text{min}}\sim(c/v_{\varphi})\Delta(T)/T$, which yields $\kappa^{(2)}\sim A_{N}/(cv_{\varphi}^{4})\gamma T\ln(v_{\varphi}^{2}T/(c^{2}\gamma))$ (see End Matter), which signals an enhanced contribution due to phonons moving in the $z$ direction. In this regime, the typical phonon momentum is large compared to the typical $\varphi$ frequencies, and unlike the previous lower-temperature regime, the phonons are now sensitive to diffusive character of $\varphi$. Such a regime is unique to phonons: for a bosonic mode of electronic origin with $c\sim v_{\varphi}$, it would only emerge at $T\sim\gamma$, an electronic scale. Here, it appears at much lower temperatures due to $c\ll v_{\varphi}$.
@@
It is useful to make a few remarks about the underlying dimensionality of the problem. The lowest temperature regimes (i), (ii) exhibiting $\kappa^{(2)}\sim T^{5}$ and $\kappa^{(2)}\sim T^{3}$ depend on the phonon dimensionality in a straightforward fashion via their density of states. However, the $\kappa^{(2)}\sim T\ln(T)$ regime (iii) involves a subtle kinematic interplay between the phonon and $\varphi$. In particular, the phonon dispersion along the $z$-direction contributes more significantly compared to the other directions, leading to the $\ln$ enhancement; the latter would be absent if the phonons were purely two-dimensional. Furthermore, $\kappa^{(2)}$ in this regime also depends crucially on the assumption of completely decoupled nature of the two-dimensional electronic layers, via the simplification, $N^{i}_{{\bm{q}}}\propto{\left|{\bm{q_{\text{2D}}}}\right|}$, instead of $\left|{\bm{q}}\right|$}. Finally, let us also consider the additional perturbation due to coupling the $\varphi$ boson to out-of-plane strain, which leads to a correction, $\lambda q_{z}^{2}$ term in $|N^{i}_{{\bm{q}}}|^{2}$ with $\lambda\ll 1$. This does not lead to qualitatively different results in regimes (i) and (ii), but it can enhance the $\ln$ factor in regime (iii) to a linear factor of $v_{\varphi}^{2}T/(c^{2}\gamma)$ (see End Matter). Since this regime is upper bounded by $T\sim\omega_{D}$, this term remains a small perturbation if $\lambda\ll\gamma c^{2}/(\omega_{D}v_{\varphi}^{2})$.
@@
Energy-relaxation due to non-linear coupling between phonon and collective mode.- Finally, we consider the contribution due to ${\mathcal{L}}_{ep}^{(3)}$ in Eq. (5), with a phonon self-energy shown in Fig. 1(d), which involves scattering between $\varphi^{2}$ and the acoustic phonons. The general expression for the flux $\kappa^{(3)}$ reads
@@
Once again, there is a complex sequence of temperature dependent crossovers that can be obtained with the following approximations: replace $\sum_{i}|L^{i}_{{\bm{k}}+{\bm{q}}/2,{\bm{q}}}|^{2}$ by $A_{L}{\left|{\bm{q_{\text{2D}}}}\right|}^{2}$, and for $T\lesssim{\omega_{\rm D}}$, include contributions from phase-space where $c{\left|{\bm{q}}\right|}\sim\nu\sim T$. We obtain
[…]
@@
[…]
@@
Connection to Experiments.- Recent optical experiments involving the technique of two-dimensional coherent spectroscopy [98, 99, 100] have studied the nonlinear response of correlated metals and its relationship to energy relaxation [101, 56], providing a further impetus to a careful theoretical examination of the questions raised above. Over a broad range of hole-dopings in a specific cuprate family [102, 103], recent pump-probe measurements [56] have identified energy relaxation as the dominant nonlinear process that governs the experimental phenomenology. At low temperatures across most of the doping phase diagram, $\Gamma_{E}\sim k_{B}T/\hbar$, with a rate that is significantly slower (approximately by an order of magnitude) compared to the Planckian momentum-relaxation rate [1, 104]. With increasing temperature, there is a tendency for $\Gamma_{E}$ to saturate across a doping dependent crossover scale. Both of these regimes are in stark contrast to the results expected in a conventional Fermi liquid coupled to acoustic phonons [35]. In THz experiments, it is the second rate $\Gamma_{ep}$ that is measured, which leads to a fast initial decay of the pump-probe signal to a plateau at the order of pico seconds, whereas the phonon-environment relaxation $\Gamma_{\text{p-env}}$ leads to a slower decay on longer timescales.
@@
To make contact with experiments, we convert $\kappa=[\kappa^{(1)}+\kappa^{(2)}+\kappa^{(3)}]$ to the energy relaxation rate, $\Gamma_{E}$, via Eq. (1). Here, due to the comparison between the heat capacities $C_{p}\sim(T/{\omega_{\rm D}})^{3}$ and $C_{e}\sim(T/E_{F})\ln(\gamma/T)$, an additional crossover is present, whose temperature scale $T_{*}\sim\sqrt{{\omega_{\rm D}}^{2}/E_{F}}$ is set by equating $C_{p}$ and $C_{e}$. Eq. (1) states that the sector with the smaller heat capacity dominates as its temperature changes more rapidly. When $T\ll T_{*}$, we have $\Gamma_{E}\approx\kappa/C_{p}$ and when $T\gg T_{*}$, $\Gamma_{E}\approx\kappa/C_{e}$. From thermodynamic measurements in a representative family of the cuprates [37], $T_{*}$ is at the order of 11-1010 K. For the experimental range of $T\sim O(100\textrm{ K})$, the latter is more relevant, so we would then divide the scaling of $\kappa$ by a power of $T$ to get scaling of $\Gamma_{E}$. Our theoretical results for $\kappa^{(3)}$ provide an appealing scenario for the experimental observations. In the experimental temperature range, the energy relaxation rate $\Gamma_{E}$ is related to $\kappa^{(3)}$ by $\Gamma_{E}\approx\kappa^{(3)}/C_{e}$. Therefore, in regime (iii) and regime (iv) we would obtain a crossover from $T\ln T$ to $T$-independent constant in $\Gamma_{E}$, which qualitatively agrees with the observation in [56] for samples near optimal doping (see Fig. 6c in [56]).
Discussion and Outlook.- We have shown that near an electronic QCP, symmetry-allowed couplings to acoustic phonons generate rich, temperature-dependent crossovers in the energy relaxation rate. When momentum relaxation in a marginal Fermi liquid is governed by electronic processes and disorder, while energy relaxation is limited by phonon coupling, there is no a priori reason for the two relaxation rates to exhibit a similar behavior. Our analysis reveals that energy relaxation follows distinct scaling behaviors across emergent low-energy regimes, independent of momentum relaxation. We also note that (a) while the electron-phonon coupling, can in principle, modify the structure of the underlying QCP [94, 95, 105, 106], the interesting regime (iii) obtained above will not be affected, and (b) the energy relaxation process we address is not affected by potential boson drag effects [107, 108, 20, 22]; see [52] for details.
Recent experiments [109, 110, 111] have motivated proposals that phonons in strongly correlated metals may lose their quasiparticle character, exhibiting Planckian lifetimes $O(\hbar/k_{B}T)$ [107, 112, 113, 114, 108]. Despite this broadening, our results for energy flux should still apply, as the relevant phonon frequencies remain $O(k_{B}T)$. A full theoretical treatment of a non-Fermi liquid entangled with incoherent phonons is left for future work. Additionally, it would be interesting to investigate how the phonon dynamics is modified by the non-Fermi liquid through numerical approaches, such as quantum Monte-Carlo [31, 12, 80, 115, 116].
@@
In this work we have focused on the normal state, but pairing fluctuations above $T_{c}$ and Bogoliubov quasiparticles below $T_{c}$ may also influence energy relaxation through their coupling to phonons [117], an interesting direction for future study. Moreover, the complex crossovers we identify—arising from mismatched velocities and dimensionality of phonons and collective modes—suggest similar behavior may appear in sound attenuation near quantum criticality [118, 119, 120, 121, 122, 123, 51, 124], which we leave for future investigation.